OpenStudy (anonymous):

(6t-5)=2(t+1)+2 help me get the solution

OpenStudy (anonymous):

2.25 6t-5=2(t=1) 6t-5=2t+2+2 6t-5=2t+4 +5 +5 6t=2t+9 -2t 4t=9 2.25